The management of a farm equipment manufacturer is evaluating a
proposal to build a plant that will manufacture lightweight tractors.
Based on a $10,000 feasibility study, the management prepared the
following cash flow projections.

The plant will be built on a plot of land that the company owns
and that currently is estimated to be worth $10 mil.

The capital expenditure at time 0 is $170 mil. The relevant CCA
rate for capital expenditure is 10%.

The new plant would produce revenues of $100 mil per year for
10 years. The project ends after 10 years and the salvage value
at the end of the project is $100 mil.

Can someone please explain how to get the number 14.29? ThanksExpansion Problem 0 ... 1 100 2 100 9 100 10 100 (10) (35) 9898 (10) 45 29.25 (10) (35) (10) 45 29.25 (10) 45 29.25 (10) 45 2
Answer #1
Year 0 1 2 3 4 5 6 7 8 9 10
CF -183 27.25 29.25 29.25 29.25 29.25 29.25 29.25 29.25 29.25 134.25
PV(CF) Formula -183 =27.25/(1+12%)^1 =29.25/(1+12%)^2 =29.25/(1+12%)^3 =29.25/(1+12%)^4 =29.25/(1+12%)^5 =29.25/(1+12%)^6 =29.25/(1+12%)^7 =29.25/(1+12%)^8 =29.25/(1+12%)^9 =134.25/(1+12%)^10
PV(CF) -183 24.33 23.32 20.82 18.59 16.60 14.82 13.23 11.81 10.55 43.22
Total PV Formula = -183+24.33+23.32+20.82+18.59+16.6+14.82+13.23+11.81+10.55+43.22
Total PV 14.29